1) Find the Taylor polynomial of order n for f(x) = ln 2x about x = 1?

2) Find the Taylor polynomial of order 3 for f(x) = 2√x about a = 4?

3)The area bounded by the curves x = 4y^2 , the line y = 1 and the y-axis in the first quadrant is rotated around the line x = 4: Determine the volume of resulting solid of revolution.


1
Expert's answer
2020-09-29T15:21:35-0400

1). We remind that the Taylor series expansion for the ln y around 1 (http://www.math.com/tables/expansion/log.htm):

"ln(y)=\sum_{n=1}^{+\infty}\frac{(-1)^{n-1}(y-1)}{n}"

After substitution "y=2x" we receive:

"ln(2x)=\sum_{n=1}^{+\infty}\frac{(-1)^{n-1}(2x-1)}{n}"

2). We remind the Taylor formula:

"f(x)=f(a)+f'(a)(x-a)+\frac{f''(a)(x-a)^2}{2!}+...+\frac{f^{(k)}(a)}{k!}(x-a)^{(k)}+h_k(x)(x-a)^{(k)},"

where "h_k(x)\rightarrow0,x\rightarrow a" .

We compute the derivatives of "f(x)=2\sqrt{x}" and get:

"f'(x)=\frac{1}{\sqrt{x}},\qquad f''(x)=-\frac{1}{2}\frac{1}{x\sqrt{x}},\qquad f'''(x)=\frac{3}{4}\frac{1}{x^2\sqrt{x}}"

Setting "k=3" and "a=4" we get:

"2\sqrt{x}=4+\frac{1}{2}(x-4)-\frac{1}{32}(x-4)^2+\frac{1}{256}(x-4)^3+h_3(x)(x-4)^3" ,

"h_3(x)\rightarrow0, x\rightarrow4"

3). The volume of the body, which is obtained after rotation of the graph is:

"\int_{0}^1\pi (4-4y^2)^2dy=\int_{0}^116\pi (1-2y^2+y^4)dy=16\pi(y-\frac{2}{3}y^3+\frac{y^5}{5})|_0^1=8\pi"

Answer: 1). "ln(2x)=\sum_{n=1}^{+\infty}\frac{(-1)^{n-1}(2x-1)}{n}"

2). "2\sqrt{x}=4+\frac{1}{2}(x-4)-\frac{1}{32}(x-4)^2+\frac{1}{256}(x-4)^3+h_3(x)(x-4)^3" ,

"h_3(x)\rightarrow0, x\rightarrow4"

3). "8\pi"